John and Peter can clean their office working together in 3 hours. working alone, it takes peter 2 hours more than it takes john working alone. how long can john do the work by himself?
Answer by nyc_function(2733)

1/3 = John and Peter working together
1/J + 2 = Peter working alone
1/J = John working alone, where J = John
1/(J + 2) + 1/J = 1/3
Can you take it from here?